# Hermitian and non-Hermitian topology in active matter

### v1

@@ 3.2.3 Higher-order topological insulator
The notion of the second-order topological insulator can be further extended to higher dimensions and orders. In three dimensions, gapless hinge modes (second order) and corner modes (third order) can appear. In particular, topological hinge states have been observed in condensed matter such as bismuth [156]. Higher-order topology is also feasible in various classical systems as we will discuss later.
3.3 Non-Hermitian topology
3.3.1 Non-Hermitian system
@@
Non-Hermitian systems, namely, nonequilibrium open systems effectively described by non-Hermitian operators [27], can exhibit a variety of phenomena without Hermitian counterparts. One of the origins of non-Hermiticity is gain and/or loss through exchanging energies with the external environment. Non-Hermitian operators can in general be nondiagonalizable, and the corresponding eigenvectors coalesce at a degenerate point of the energy eigenvalues, which is known as an exceptional point [157, 158, 159]. An exceptional point can be interpreted as a topological object in gapless non-Hermitian systems, and its topological nature can lead to anomalous bulk-edge correspondence as explained below. Another possible origin of non-Hermiticity is nonreciprocity caused by, e.g., the unidirectional flow of energies and particles, which can lead to numerous localized modes. This localization phenomenon, also known as the non-Hermitian skin effect [160, 161, 162], can be attributed to nontrivial non-Hermitian topology originating from a unique complex energy gap. In the following sections, we shall explain the unconventional bulk-edge correspondence between a topological invariant defined by complex energy bands and the aggregation of the bulk modes at the boundaries. As such, studies of non-Hermitian topology might deepen our understanding of the robust behavior in nonequilibrium open systems including active matter.
@@
Non-Hermitian physics has been experimentally studied in optics and photonics in controllable settings [163, 164, 165, 166, 167]. More recently, non-Hermitian systems have been also explored in the fields of atomic physics [168] and magnonics [169, 170]. Meanwhile, various classical systems, such as acoustics [132, 133, 171, 172], can also emulate non-Hermitian models to realize and observe various types of non-Hermitian phenomena. In recent years, the concept of non-Hermitian topology has also found applications to active matter; owing to its inherently nonequilibrium nature, active matter can naturally be described as non-Hermitian systems as detailed below. For instance, unidirectional flow induced by self-propelled force can lead to nonreciprocity, which makes the generator of the time evolution non-Hermitian.
3.3.2 Point gap and line gap
@@
Figure 9: Schematic figures of energy gaps. (a) Energy gap in Hermitian systems. (b) Point gap in non-Hermitian systems. The base energy $E_{\rm B}$ is located inside the energy band on the complex plane. (c) Line gap in non-Hermitian systems. The red line divides the two sets of the energy bands.
@@
The key ingredient of defining topological invariants is an energy gap, the range in which no energy eigenvalues exist. There is no ambiguity in the definition of the energy gap for Hermitian systems [Fig. 9(a)] because the energy eigenvalues are always real-valued, for which the corresponding topological invariant is well-defined as explained in the previous sections. Meanwhile, in non-Hermitian systems, the fact that the energy eigenvalues can take complex values leads to several possible definitions of an energy gap [27]. There are two common ways to do so; first, non-Hermitian systems are said to have a point gap when the energy eigenvalues open a gap at the base energy on the complex plane. As shown in Fig. 9(b), when the energy band forms a closed curve, the point gap opens inside and outside the curve, closing on the curve [173]. Accordingly, the topological invariant related to the point gap changes its values depending on the base energy, as discussed below. The other possible definition is a line gap where two sets of energy eigenvalues are separated by a line determining the gap on the complex plane [Fig. 9(c)] [174]. The topology associated with the line gap is essentially equivalent to the Hermitian topology because any non-Hermitian Hamiltonian can be continuously deformed to a Hermitian Hamiltonian while keeping the line gap if any [27, 174]. The line-gap topology can give rise to localization modes with discrete energy eigenvalues similar to Hermitian cases.
@@ 3.3.3 Non-Hermitian skin effect in one-dimensional systems
Figure 10: Energy eigenvalues and spatial profiles of the bulk eigenstates of the Hatano-Nelson model. (a) Energy band obtained from Eq. (46) (blue) and energy eigenvalues of the finite-size system with open boundary conditions (red). (b) Localization of the bulk eigenstates. The system parameters are set to be $t_{L}=0.3$ and $t_{R}=1$.
We first illustrate the notion of non-Hermitian topology on the basis of the bulk-edge correspondence in one-dimensional (1D) non-Hermitian tight-binding systems. The key idea is that the winding of the complex-valued energy band can characterize the point-gap topology [175, 176]. Specifically, for a non-Hermitian Bloch Hamiltonian denoted by $H\left(k\right)$, the topological invariant detecting the point-gap topology is defined as
[…]
which is called the energy winding number. The base energy $E_{\rm B}$ can be any one point on the complex plane. To see a nonzero energy winding number, we consider the Hatano-Nelson model [177], a prototypical tight-binding model with a nontrivial non-Hermitian topology. Its real-space Hamiltonian reads
@@
where $t_{L}$ and $t_{R}$ are real-valued parameters, which express asymmetric hopping amplitudes. We show the energy band of the Hatano-Nelson model in the case of $t_{L}<t_{R}$ in Fig. 10(a). From Eq. (44), one can immediately check that the energy winding number takes the value of either 00 or $-1$ depending on whether the base energy is located outside or inside the energy band on the complex plane. Namely, the energy winding number changes when the point gap closes.
We next explain the physical meaning of the energy winding number. For a non-Hermitian Hamiltonian ${\cal H}$, we can define the extended Hermitian Hamiltonian as
~$\tilde{\cal H}=\left(\begin{array}[]{cc}O&{\cal H}-E\\ {\cal H}^{\dagger}-E^{\ast}&O\end{array}\right),$ (49)
@@
where $\Gamma={\rm diag}\left(1,-1\right)$. Remarkably, the classification of the point-gap topology of the non-Hermitian Hamiltonian is equivalent to the topological classification of the extended Hermitian Hamiltonian with the chiral symmetry [173]. For example, since the Hatano-Nelson model has no symmetries, the extended Hermitian Hamiltonian obtained from Eq. (45) belongs to Class AIII in the conventional AZ symmetry class. The previous section shows that the Hermitian topology in Class AIII can be characterized by a $Z$ topological invariant. Accordingly, the point-gap topology of the Hatano-Nelson model can also be characterized by a $Z$ topological invariant, which is nothing but the energy winding number.
The previous section shows that the Hermitian topology allows us to establish the bulk-edge correspondence between a topological invariant defined in bulk and the existence of spatially localized in-gap states. The topological equivalence between non-Hermitian systems and extended Hermitian systems above indicates that a similar bulk-edge correspondence might be also established in non-Hermitian systems. Indeed, one can show that the nonzero energy winding number corresponds to the localization of the bulk eigenstates to edges [175, 176]. Such an unconventional localization phenomenon is called the non-Hermitian skin effect [160]. For example, in the Hatano-Nelson model, all the eigenstates are localized at one edge [Fig. 10(b)] due to the nonzero winding number. The localization can be understood as the accumulation of the eigenstates under unidirectional flow owing to the asymmetric hopping amplitudes. We note that the corresponding energy eigenvalues in a finite-size system with open boundary conditions take real values, which are distinct from the energy band obtained from the Bloch Hamiltonian (46). So far, many physical implementations of the non-Hermitian skin effect have been reported in various areas of research, such as photonics [178, 179, 180, 181, 182], ultracold atoms [183], acoustics [184], mechanics [9, 185, 186], and electrical circuits [187, 188].
@@
We note that this operation is not equivalent to the conventional time-reversal symmetry because of $H^{\rm T}\neq H^{\ast}$ in non-Hermitian systems. This symmetry acts as an inverter of the spin direction corresponding to the self-rotation of particles, leading to the Kramers degeneracy. The corresponding extended Hermitian Hamiltonian then belongs to Class D, and it is classified by a $Z_{2}$ topological invariant. The physical meaning of this classification is that odd numbers of Kramers pairs appear as topological edge states. Accordingly, the $Z_{2}$ point-gap topology corresponds to the unconventional non-Hermitian skin effect, where the bulk eigenstates with spin-up are localized at one end, and those with spin-down are localized at the other, in contrast to the localization nature induced by the $Z$ point-gap topology. Thus, the symmetry-protected non-Hermitian skin effect can be understood as an analog of a quantum spin Hall insulator.
The point-gap topology plays a crucial role also in stochastic processes, which can be described by the master equation given as follows:
@@
where ${\bm{P}}$ and $W$ represent a vector of the occupation probabilities of the states and the transition matrix, respectively. We consider a 1D Markov system with spatial translation symmetry, in which the transition matrix can be diagonalized by employing the representation $W\left(k\right)$ in the Fourier basis. One can then extract the point-gap topology of the transition matrix. For example, the transition matrix akin to the Hatano-Nelson model has a nonzero winding number. This simply means that the random walker described by the transition matrix flows in one direction due to the asymmetric hopping rate and eventually reaches the steady state which is localized at the edge. Indeed, Ref. [190] has established the correspondence of the nonzero topological invariant and the existence of the localized steady state. We will give a specific definition of the winding number and a qualitative description of the localized steady state later. Thus, the point-gap topology allows us to predict the nature of localization in stochastic processes.
@@ 3.3.4 Non-Hermitian skin effect in two-dimensional systems
Figure 11: Geometry-dependent skin effect of the model (54). (a) Energy band. (b) Spatial profile of the bulk eigenstates on the diamond geometry. The bulk eigenstates are localized at the edges.
@@
We next review the non-Hermitian topology and the associated non-Hermitian skin effect in two-dimensional (2D) non-Hermitian tight-binding systems. We first consider the 2D square-lattice system composed of horizontal and vertical Hatano-Nelson chains. One can immediately infer that this system can be characterized by two energy winding numbers defined in Eq. (44). Indeed, the corresponding non-Hermitian skin effect localizes all the bulk eigenstates at one corner [191]. The localization can be readily understood as the flow of the eigenstates toward the corner due to the asymmetric hopping amplitudes in all directions. Such localization of the bulk eigenstates and its manipulation have been experimentally realized in acoustics [192] and photonics [193]. However, the system does not exhibit phenomena unique to 2D systems in the sense that the observed non-Hermitian skin effect merely inherits 1D nature.
@@
Recent works have pointed out that the point-gap topology corresponds not only to the Hermitian first-order topology but also to the second-order topology [194, 195, 196]. Specifically, one can define topological invariants of non-Hermitian systems via the extended Hermitian Hamiltonian which exhibits second-order topological phases. When the topological invariant takes a nonzero value, several eigenstates are then localized at the corners. This phenomenon is called the second-order topological non-Hermitian skin effect, and it is unique to 2D systems since the localization originates from second-order topology which can appear in more than 2D systems. Remarkably, it is possible to realize the second-order non-Hermitian skin effect without nonreciprocal interaction or unidirectional flow in contrast to 1D non-Hermitian systems mentioned above.
@@
The additional mechanism to localize the eigenstate has been proposed by combining the point-gap topology and the first-order Hermitian topology. This type of localization is called the hybrid skin-topological effect. The key idea is to manipulate the topological edge states so that the winding number (44) takes nonzero values. For example, in the Chern insulator, adding gain and loss to the bulk [197, 198, 199] or the edges [200] induces the topological edge states localized at the corners. Additionally, Ref. [195] has proposed a model constructed by stacking the Hatano-Nelson model so that the net flow becomes zero and by introducing the SSH-model-like hopping amplitudes along the stacking direction. As a result, the non-Hermitian skin effect localizes the topological edge states induced by the SSH-model-like hopping amplitudes at the corners. Said differently, the model construction gives rise to hybridizing the topological edge states from the SSH model and the skin states from the Hatano-Nelson model. The hybrid skin-topological effect has been experimentally observed in electrical circuits [201] and active matter [202] as we will review in Sec. 4.3.1.
@@
Finally, we briefly review a novel non-Hermitian skin effect found in recent years. We consider a one-band model described by a Bloch Hamiltonian $H\left({\bm{k}}\right)$ and assume that all the hopping amplitudes of the model exhibit reciprocity. Reference [161] has proposed that, for any ${\bm{k}}_{\rm B}$ in the Brillouin zone, one can define the winding number as follows:
[…]
@@
where $\Gamma_{{\bm{k}}_{\rm B}}$ represents the infinitesimal counterclockwise loop enclosing ${\bm{k}}_{r}$, and $E_{\rm B}=H\left({\bm{k}}_{\rm B}\right)$. The notable point is that the energy band has a finite area on the complex plane when Eq. (53) equals nonzero. The existence of the spectral area indicates that one can take an integral path inside the area so that the contour integral defined in Eq. (44) can take nonzero values, which leads to the non-Hermitian skin effect. Said differently, the finite spectral area gives rise to the non-Hermitian skin effect even though the system does not have nonreciprocal hopping amplitudes. The key point is that the occurrence of the non-Hermitian skin effect depends on the shape of the system.
@@
One can immediately confirm that the winding number is $1$ because the energy band has a finite area [Fig. 11(a)]. The bulk eigenstates are then localized at the edges of the diamond geometry [Fig. 11(b)], and the localization states are qualitatively distinct from the corner-localized states introduced above. We remark that the non-Hermitian skin effect does not occur on the square geometry. This phenomenon is called the geometry-dependent skin effect, which has been experimentally observed in phononics [203].

### v2

@@ 3.2.3 Higher-order topological insulator
The notion of the second-order topological insulator can be further extended to higher dimensions and orders. In three dimensions, gapless hinge modes (second order) and corner modes (third order) can appear. In particular, topological hinge states have been observed in condensed matter such as bismuth [161]. Higher-order topology is also feasible in various classical systems as we will discuss later (cf. Sec. 5.3.1).
4 Non-Hermitian band topology
4.1 Non-Hermitian system
@@
Non-Hermitian systems, namely, nonequilibrium open systems effectively described by non-Hermitian operators [25], can exhibit a variety of phenomena without Hermitian counterparts. One of the origins of non-Hermiticity is gain and/or loss through exchanging energies with the external environment. Non-Hermitian operators can in general be nondiagonalizable, and the corresponding eigenvectors coalesce at a degenerate point of the energy eigenvalues, which is known as an exceptional point [162, 163, 164]. An exceptional point can be interpreted as a topological object in gapless non-Hermitian systems, and its topological nature can lead to anomalous bulk-edge correspondence as explained below. Another possible origin of non-Hermiticity is nonreciprocity caused by, e.g., the unidirectional flow of energies and particles, which can lead to numerous localized modes. This localization phenomenon, also known as a non-Hermitian skin effect [165, 166, 167, 168], can be attributed to nontrivial non-Hermitian topology originating from a unique complex-valued energy gap. In the following sections, we shall explain the unconventional bulk-edge correspondence between a topological invariant defined by complex-valued energy bands and the aggregation of the bulk modes at the boundary. As such, studies of non-Hermitian topology might deepen our understanding of the robust behavior in nonequilibrium open systems including active matter.
@@
Non-Hermitian physics has been experimentally studied in optics and photonics in controllable settings [169, 170, 171, 172, 173]. More recently, non-Hermitian systems have been also explored in the fields of atomic physics [174] and magnonics [175, 176]. Meanwhile, various classical systems, such as acoustics [135, 136, 177, 178], can also emulate non-Hermitian models to realize and observe various types of non-Hermitian phenomena. In recent years, the concept of non-Hermitian topology has also found applications to active matter; owing to its inherently nonequilibrium nature, active matter can naturally be described as non-Hermitian systems as detailed below (cf. Sec. 5.3). For instance, unidirectional flow induced by self-propelled force can lead to nonreciprocity, which gives rise to non-Hermitian dynamics.
4.2 Point gap and line gap
@@
Figure 9: Schematic figures of energy gaps. (a) Energy gap in Hermitian systems. (b) Point gap in non-Hermitian systems. The base energy $E_{\rm B}$ is located inside the energy band on the complex-valued energy plane. (c) Line gap in non-Hermitian systems. The red line divides the two sets of the energy bands.
@@
The key ingredient of defining topological invariants is an energy gap, the range in which no energy eigenvalues exist. There is no ambiguity in the definition of the energy gap for Hermitian systems [Fig. 9(a)] because the energy eigenvalue is always real-valued, for which the corresponding topological invariant is well-defined as explained in Sec. 3.2. Meanwhile, in non-Hermitian systems, the fact that the energy eigenvalue can take complex values leads to several possible definitions of energy gaps [25]. There are two common ways to do so; first, non-Hermitian systems are said to have a point gap when the energy eigenvalue opens a gap at the base energy on the complex-valued energy plane. As shown in Fig. 9(b), when the energy band forms a closed curve, the point gap opens inside and outside the curve, closing on the curve [179]. Accordingly, the topological invariant related to the point gap changes its values depending on the base energy, as discussed below. The other possible definition is a line gap where two sets of energy eigenvalues are separated by a line determining the gap on the complex-valued energy plane [Fig. 9(c)] [180]. The topology associated with the line gap is essentially equivalent to the Hermitian topology because any non-Hermitian Hamiltonian can be continuously deformed to a Hermitian Hamiltonian while keeping the line gap if any [25, 180]. The line-gap topology can give rise to localization modes with discrete energy eigenvalues similar to Hermitian cases.
4.3 Non-Hermitian skin effect
@@ 4.3.1 Non-Hermitian skin effect in one-dimensional systems
Figure 10: Energy eigenvalues and spatial profile of the bulk eigenstates of the Hatano-Nelson model. (a) Energy band obtained from Eq. (47) (blue) and energy eigenvalues of the finite-size system with an open boundary condition (red). (b) Localization of the bulk eigenstates. The system parameters are set to be $t_{L}=0.3$ and $t_{R}=1$.
We first illustrate the notion of non-Hermitian topology on the basis of the bulk-edge correspondence in one-dimensional (1D) non-Hermitian tight-binding systems. The key idea is that the winding of the complex-valued energy band can characterize the point-gap topology [181, 182, 183]. Specifically, for a non-Hermitian Bloch Hamiltonian denoted by $H\left(k\right)$, the topological invariant detecting the point-gap topology is defined as
$W_{\rm P}=\frac{1}{2\pi i}\int_{0}^{2\pi}dk\frac{d}{dk}\log\det\left[H\left(k\right)-E_{B}\right],$ (45)
which is called an energy winding number. The base energy $E_{\rm B}$ can be any one point on the complex-valued energy plane. To see a nonzero energy winding number, we consider the Hatano-Nelson model [184], a prototypical tight-binding model with a nontrivial non-Hermitian topology. Its real-space Hamiltonian reads
@@
where $t_{L}$ and $t_{R}$ are real-valued parameters, which express asymmetric hopping amplitudes. We show the energy band of the Hatano-Nelson model in the case of $t_{L}<t_{R}$ in Fig. 10(a). From Eq. (45), one can immediately check that the energy winding number takes the value of either 0 or 1-1 depending on whether the base energy is located outside or inside the energy band. Namely, the energy winding number changes when the point gap closes.
[…]
$\tilde{\cal H}=\left(\begin{array}[]{cc}O&{\cal H}-E\\ {\cal H}^{\dagger}-E^{\ast}&O\end{array}\right),$ (50)
@@
where $\Gamma$ is defined in the form of Eq. (42). Remarkably, the classification of the point-gap topology of the non-Hermitian Hamiltonian is equivalent to the topological classification of the extended Hermitian Hamiltonian with the chiral symmetry [179]. For example, since the Hatano-Nelson model has no symmetries, the extended Hermitian Hamiltonian obtained from Eq. (46) belongs to Class AIII in the conventional AZ symmetry class. The Hermitian topology in Class AIII can be characterized by a $Z$ topological invariant, as discussed in the SSH model (cf. Sec. 3.2.2). Accordingly, the point-gap topology of the Hatano-Nelson model can also be characterized by a $Z$ topological invariant, which is nothing but the energy winding number.
Section 3.2 shows that the Hermitian topology allows us to establish the bulk-edge correspondence between a topological invariant defined in bulk and the existence of spatially localized in-gap states. The topological equivalence between non-Hermitian systems and extended Hermitian systems above indicates that a similar bulk-edge correspondence might be also established in non-Hermitian systems. Indeed, one can show that the nonzero energy winding number corresponds to the localization of the bulk eigenstate to the edge [182, 183]. Such an unconventional localization phenomenon is called the non-Hermitian skin effect [165]. For example, in the Hatano-Nelson model, all the eigenstates are localized at one edge [Fig. 10(b)] due to the nonzero winding number. The localization can be understood as the accumulation of the eigenstates under unidirectional flow owing to the asymmetric hopping amplitudes. We note that the corresponding energy eigenvalues under an open boundary condition take real values, which are distinct from the energy band obtained from the Bloch Hamiltonian (47). So far, many physical implementations of the non-Hermitian skin effect have been reported in various areas of research, such as photonics [185, 186, 187, 188, 189], ultracold atoms [190], acoustics [191], mechanics [9, 192, 193], and electrical circuits [194, 195]. On another front, a recent previous work has proposed that the boundary sensitivity of the energy eigenvalues is applicable for highly sensitive sensors [196].
@@
We note that this operation is not equivalent to the conventional time-reversal symmetry because of $H^{\rm T}\neq H^{\ast}$ in non-Hermitian systems. This symmetry acts as an inverter of the spin direction corresponding to the self-rotation of particles, leading to the Kramers degeneracy. The corresponding extended Hermitian Hamiltonian then belongs to Class D, and it is classified by a $Z_{2}$ topological invariant. The physical meaning of this classification is that odd numbers of Kramers pairs appear as topological edge states. Accordingly, the $Z_{2}$ point-gap topology corresponds to the unconventional non-Hermitian skin effect, where the bulk eigenstates with spin-up are localized at one end, and those with spin-down are localized at the other, in contrast to the localization nature induced by the $Z$ point-gap topology. Thus, the symmetry-protected non-Hermitian skin effect can be understood as an analog of a quantum spin Hall insulator (cf. Sec. 3.2.2).
The point-gap topology plays a crucial role also in stochastic processes, which can be described by a master equation given as follows:
@@
where $\bm{P}$} and $W$ represent a vector of occupation probabilities of states and a transition matrix, respectively. We consider a 1D Markov system with a spatial translation symmetry, in which the transition matrix can be diagonalized by employing the representation $W\left(k\right)$ in the Fourier basis. One can then extract the point-gap topology of the transition matrix. For example, the transition matrix akin to the Hatano-Nelson model has a nonzero winding number. This simply means that a random walker described by the transition matrix flows in one direction due to an asymmetric hopping rate and eventually reaches a steady state which is localized at the edge. Indeed, Ref. [198] has established the correspondence of the nonzero topological invariant and the existence of the localized steady state. We will give a specific definition of the winding number and a qualitative description of the localized steady state in Sec. 4.3.4. Thus, the point-gap topology allows us to predict the nature of localization in stochastic processes.
@@ 4.3.2 Non-Hermitian skin effect in two-dimensional systems
Figure 11: Geometry-dependent skin effect of the model ($5$). (a) Energy band. (b) Density of state on the diamond geometry. The bulk eigenstates are localized at the edges.
@@
We next review the non-Hermitian topology and the associated non-Hermitian skin effect in two-dimensional (2D) non-Hermitian tight-binding systems. We first consider a 2D square-lattice system composed of horizontal and vertical Hatano-Nelson chains. One can immediately infer that this system can be characterized by two energy winding numbers defined in Eq. (45). Indeed, the corresponding non-Hermitian skin effect localizes all the bulk eigenstates at one corner [199]. The localization can be readily understood as the flow of the eigenstates toward the corner due to the asymmetric hopping amplitudes in all the directions. Such localization of the bulk eigenstate and its manipulation have been experimentally realized in acoustics [200] and photonics [201]. However, the system does not exhibit phenomena unique to 2D systems in the sense that the observed non-Hermitian skin effect merely inherits 1D nature.
@@
Recent works have pointed out that the point-gap topology corresponds not only to the Hermitian first-order topology but also to the second-order topology [202, 203, 204], which we have discussed in Sec. 3.2.3. Specifically, one can define a topological invariant of a non-Hermitian system via an extended Hermitian Hamiltonian which exhibits a second-order topological phase. When the topological invariant takes a nonzero value, several eigenstates are then localized at the corners. This phenomenon is called a second-order topological non-Hermitian skin effect, and it is unique to 2D systems since the localization originates from second-order topology which can appear in more than 2D systems. Remarkably, it is possible to realize the second-order non-Hermitian skin effect without nonreciprocal interaction or unidirectional flow in contrast to 1D non-Hermitian systems discussed in Sec. 4.3.1.
@@
The additional mechanism to localize eigenstates has been proposed by combining the point-gap topology and the first-order Hermitian topology. This type of localization is called a hybrid skin-topological effect. The key idea is to manipulate the topological edge state so that the winding number (45) takes a nonzero value. For example, in the Chern insulator (cf. Sec. 3.2.1), adding gain and loss to the bulk [205, 206, 207] or the edges [208] induces the topological edge state localized at the corners. Additionally, Ref. [203] has proposed a model constructed by stacking the Hatano-Nelson model so that the net flow becomes zero and by introducing the SSH-model-like hopping amplitude along the stacking direction. As a result, the non-Hermitian skin effect localizes the topological edge states induced by the SSH-model-like hopping amplitude at the corners. Said differently, the model construction gives rise to hybridizing the topological edge states from the SSH model and the skin states from the Hatano-Nelson model. The hybrid skin-topological effect has been experimentally observed in electrical circuits [209] and active matter [210] as we will review in Sec. 5.3.1.
@@
Finally, we briefly review a novel non-Hermitian skin effect found in recent years. We consider a one-band model described by a Bloch Hamiltonian $H\left({\bm{k}}\right)$ and assume that all hopping amplitudes of the model exhibit reciprocity. Reference [166] has proposed that, for any ${\bm{k}}_{\rm B}$ in the Brillouin zone, one can define the winding number as follows:
$W_{\rm 2D}=\frac{1}{2\pi i}\oint_{\Gamma_{{\bm{k}}_{\rm B}}}d{\bm{k}}\cdot{\bm{\nabla}}_{\bm{k}}\log\det\left[H\left({\bm{k}}\right)-E_{\rm B}\right],$ (54)
@@
where $\Gamma_{{\bm{k}}_{\rm B}}$ represents an infinitesimal counterclockwise loop enclosing ${\bm{k}}_{r}$, and $E_{\rm B}=H\left({\bm{k}}_{\rm B}\right)$. The notable point is that the energy band has a finite area on the complex-valued energy plane when Eq. (54) equals nonzero. The existence of the spectral area indicates that one can take an integral path inside the area so that the contour integral defined in Eq. (45) can take nonzero values, which leads to the non-Hermitian skin effect. Said differently, the finite spectral area gives rise to the non-Hermitian skin effect even though the system does not have nonreciprocal hopping amplitudes. The key point is that the occurrence of the non-Hermitian skin effect depends on the shape of the system.
@@
One can immediately confirm that the winding number (54) is $1$ in the finite area occupied by the energy band [Fig. 11(a)]. The bulk eigenstates are then localized at the edges of the diamond geometry [Fig. 11(b)], and the localization states are qualitatively distinct from the corner-localized states introduced above. We remark that the non-Hermitian skin effect does not occur on the square geometry. This phenomenon is called a geometry-dependent skin effect, which has been experimentally observed in phononics [211].